A pulley is used to lift a car engine . The engine is lifted 35.6 cm for every 1.07 m the rope is pulled . What is the IMA of the pulley


1
Expert's answer
2021-06-15T17:00:29-0400

Gives

d1=3.56m

d2=1.07m

"W=F.dr"

Win=Wout

"(F.d)_{in}=({F.d})_{out}"

Put value

"F_{in}\times3.56=F_{out}\times1.07"

"\frac{F_{out}}{F_{in}}=\frac{3.56}{1.07}=3.32"

We know that

IMA="\frac{F_{out}}{F_{in}}"

Put value

"IMA=\frac{3.56}{1.07}=3.32"